A network provider investigates the load of its network. The number of concurrent users is recorded at Seventy - seven locations (thousands of people),

55.3 55.3 55.3 55.9 55.9 55.9 55.9 56.1 56.1 56.1 56.1 56.1 56.1 56.8

56.8 57.0 57.0 57.0 57.8 57.8 57.8 57.9 57.9 57.9 58.8 58.8 58.8 59.8

59.8 59.8 62.2 62.2 63.8 63.8 63.8 63.9 63.9 63.9 64.7 64.7 64.7 65.1

65.1 65.1 65.3 65.3 65.3 65.3 67.4 67.4 67.4 67.4 68.7 68.7 68.7 68.7

69.0 70.4 70.4 71.2 71.2 71.2 73.0 73.0 73.1 73.1 74.6 74.6 74.6 74.6

79.3 79.3 79.3 79.3 83.0 83.0 83.0

1) Construct a grouped frequency distribution for the data using five classes. Add the Relative frequency and the Percent frequency tables to the main frequency table. 


1
Expert's answer
2020-08-17T12:42:13-0400

We have the sample size "n=77", the number of intervals (classes) "s=5", max = 83, min = 55.3.

We compute the length of each interval "h=\frac{max-min}{s}=\frac{83-55.3}{5}=5.54".

So we get the upper limits: 60.84, 66.38, 71.92, 77.46, 83.

Using COUNTIFS function in Excel we find the frequency (the number of observations in the intervals), the relative frequency (the number of observations in the interval/77) and percent frequency ((the number of observations in the interval/77)*100%).